raw a flowchart for converting a total of the 28 different faharenheit temperature to celsius temperature?
1
Expert's answer
2020-03-02T09:12:21-0500

As per the given question,

Temperature reading on faharenheit scale =28

We know that "C=\dfrac{5(F-32)}{9}"

"C=\dfrac{5(28-32)}{9}=\dfrac{20}{9}^\circ C"

Flow diagram is given below,